OpenStudy (anonymous):

solve e^(2x)+2e^x=24

OpenStudy (anonymous):

transform into a quadratic eeqn

OpenStudy (anonymous):

and then solve it

OpenStudy (anonymous):

how do I turn it into a quadratic?

OpenStudy (anonymous):

just put e^x = t

OpenStudy (antman):

e^(2x)+2e^x=4 let x be e^x be a variable, then y^2+2x-24=0. solve for y, answer is y=4 , but you know that y=e^x which is 4=e^x solve fr X and ln(4) or 2ln(2)

OpenStudy (anonymous):

omg thank you so much!